# Generalizable Graph Neural Networks for Robust Power Grid Topology Control

### v2

@@ Abstract
The energy transition necessitates new congestion management methods. One such method is controlling the grid topology with machine learning (ML). This approach has gained popularity following the Learning to Run a Power Network (L2RPN) competitions. Graph neural networks (GNNs) are a class of ML models that reflect graph structure in their computation, which makes them suitable for power grid modeling. Various GNN approaches for topology control have thus been proposed. We propose the first GNN model for grid topology control that uses only GNN layers. Additionally, we identify the busbar information asymmetry problem that the popular homogeneous graph representation suffers from, and propose a heterogeneous graph representation to resolve it. We train both homogeneous and heterogeneous GNNs and fully connected neural networks (FCNN) baselines on an imitation learning task. We evaluate the models according to their classification accuracy and grid operation ability. We find that the heterogeneous GNNs perform best on in-distribution networks, followed by the FCNNs, and lastly, the homogeneous GNNs. We also find that both GNN types generalize better to out-of-distribution networks than FCNNs.
@@
Heterogeneous over homogeneous graphs increase graph neural network expressiveness
Heterogeneous graph neural networks are the highest-performing models
Graph neural networks generalize better to out-of-distribution networks
Machine learning agents obtain good performance at high efficiency
@@ 1 Introduction
The energy transition is crucial for ensuring society’s sustainability and future security. Transmission system operators (TSOs) play a crucial role in this transition. They are consequently facing new operational challenges [9]. Grid congestion is a problem that TSOs are already encountering but which the energy transition will exacerbate [42]. If not mitigated, grid congestion could make the future availability of energy unreliable.
There are several actions that TSOs can take to alleviate grid congestion [17]. Presently, redispatching and curtailment are the primary methods. However, these actions are often costly, disruptive, or not applicable. Topology control presents an opportunity for remedying grid congestion without these problems [41]. Congestion can be relieved by exploiting the flexibility in the network topology. However, this technique mostly remained unutilized, as the space of potential topologies is combinatorially large.
Due to the recent congestion management challenges, there has been a new interest in topology control. This interest led to the development of the GridOptions tool for remedial action recommendation by Viebahn et al. [41]. This tool validates the practical value of topology control. Even simple topological strategies could reduce line congestion by ten to twenty percent. The recommended strategies improved upon strategies without topological actions or those considered by grid operators. However, while the results are promising, further developments are necessary for broader applicability. The tool was only applied to a small subset of the Dutch power grid, featuring nine substations. Additionally, the inference time of the tool is relatively long: one hour. The next challenges involve scaling this approach and reducing inference time.
Machine learning (ML) presents a solution to these challenges [42]. ML models can learn which topologies are suitable in which situations and predict these topologies much quicker than traditional computational methods. To facilitate research in this area, the Learning to Run a Power Network (L2RPN) competitions were hosted [28, 27]. Furthermore, the Grid2Op Python library was developed. [^1] https://grid2op.readthedocs.io/ This library greatly simplified the development and evaluation of ML methods for grid topology control. These developments kick-started research into ML methods for grid topology control.
Classical ML models are typically designed for tabular data and hence fail to capture the graph structure of power grids. This is problematic as the grid topology frequently changes, e.g., because of maintainance. Power grid operation requires adapting to such changes. Graph neural networks (GNNs) [36] are a class of ML models for graph-structured data. The computation in GNNs can reflect arbitrary graphs, which allows them to exploit a graph’s structure. This makes GNNs suitable for modeling power grids. GNNs have thus been applied to many grid operation tasks [5, 4, 31, 11, 18, 13, 30].
Various GNN approaches have been proposed for grid topology control [45, 43, 44, 35, 40, 33]. This research has demonstrated the feasibility of GNN approaches for grid topology control. However, important topics have not yet been addressed:
@@
    All approaches that directly predict topology actions with GNNs only use GNN layers as feature extractors or embedders [45, 43, 44, 33]. Fully-connected or attention layers follow the GNN layers. No research has investigated ’fully graphical’ neural networks for grid topology control. Purely GNN approaches might be able to exploit the topology more deeply.
@@
    All present literature uses the graph representation specified by which busbars and power lines objects share. However, this is only one of many ways to represent the power grid as a graph. There has not been any study that compares graph representations and their effect on GNN performance.
  3. 3.
    Grid changes are frequent, and operation needs to continue despite these changes. For robust operation, models must generalize to potentially unexpected network states. A previous competition has focused on generalization to out-of-distribution injection profiles [27]. However, no work has investigated generalization to out-of-distribution networks.
This study aims to fill the aforementioned gaps in applying GNNs to grid topology control. We develop fully GNN models that use a richer graph representation. The main contributions of this study are:
@@
    We propose the first fully graphical neural network approach for grid topology control. The models are trained on a multi-label binary classification imitation learning task involving various networks and topologies. We compare the GNNs against fully connected neural networks (FCNNs). We evaluate the models by their classification accuracy and ability to operate the power grid as agents. We also consider expert and hybrid agents and compare their inference speed.
@@
    We first analyze the effect of grid graph representation on GNN performance. The graph representation in the literature models current but not potential object-busbar attachments. This results in busbar information asymmetry, which blocks information flow and impairs GNN expressiveness. We propose a heterogeneous graph representation that models current and potential connections, which remedies busbar information asymmetry. We evaluate both graph representations with GNNs.
  3. 3.
    We first investigate the ability of graph neural networks to generalize topology control to out-of-distribution grids. Generalization is evaluated w.r.t. model accuracy and the ability of agents to operate the power grid, and contrasted against FCNNs.
This paper is organized as follows. Section 2 describes related work in ML grid topology control. Section 3 describes the power grid setup used in this study. Section 4 describes the methods used to develop the ML models. Section 5 describes the evaluation and analysis of the ML models. Sections 6 and 7 provide discussion and recommendations for future work, respectively. The code for this project is available. [^2] https://github.com/MatthijsdeJ/GNN_PN_Imitation_Learning Table 1 lists the meaning of mathematical symbols used in this paper. Table 2 lists the glossary.
@@
$O$ The set of grid objects (i.e., generators, loads, and line endpoints).
$S$ The set of substations.
$\textbf{x}_{u}$ The features of a GNN model at node $u\in O$.
$\textbf{h}_{u,k}$ The node embeddings of a GNN model at node $u\in O$ in layer $k$.
$\mathcal{N}(u)$ The neighbors of node $u\in O$.
$W$ A weight matrix.
$\sigma$ An activation function.
$\textbf{p}=(0,1)^{|O|}$ The output of a ML model.
$\textbf{y}=\{0,1\}^{|O|}$ The target of a ML model.
$\alpha=0.1$ The label weight hyperparameter.
$\textbf{w}=\{\alpha,1\}^{|O|}$ The label weights for the loss.
@@
$\{\textbf{p}_{s}\}{}\forall s\in S$ A partition of the output vector p into subvectors $\textbf{p}_{s}$. Subvector $\textbf{p}_{s}$ corresponds to substation $s\in S$.
$\eta$ = 0.97 The activity threshold parameter.
$\theta=1.0$ The risk threshold parameter, used by various agents.
Table 2: Glossary
Term Meaning
Busbar Information Asymmetry An asymmetry in the availability of information of objects on busbars.
Full-network Regime An environmental regime without statically or randomly disabled lines.
In-distribution Data (ID) Data with networks that the models are exposed to during training.
Injection The generators and loads of the environment.
Loading The ratio of a line’s current over its thermal limit.
Multi-label Binary Classification A learning task where the target consists of multiple binary values.
N-1 Network A variation of the default network with a single line disabled.
N-1 Redundancy The state of the network being resilient to the disablement of any single line.
Object Generators, loads, and line endpoints in the Grid2Op environment.
Out-of-distribution Data (OOD) Data with networks that the models are not exposed to during training; used to investigate generalization.
Planned-outage Regime An environmental regime with statically disabled lines.
Topology Reversal The periodical resetting of the topology to the default topology.
Topology Vector The Grid2Op variable that specifies the object-busbar configurations.
Unplanned-outage Regime An environmental regime with an opponent that randomly disables lines.
2 Related Work
The L2RPN competitions piqued the academic interest in grid topology control with machine learning. Prior research into topology control used linear programming [10, 49], but such methods are too rigid and/or slow to be practical. The first L2RPN competition premiered in 2019. It focused on operating scenarios of the EEE14-bus network within a specific time frame [28]. The following L2RPN competition was featured at WCCI in 2020 and featured the larger IEEE118-bus network [29]. The next competition at NeurIPS 2020 [27] introduced a robustness track with unplanned outages and an adaptability track with out-of-distribution injection profiles. Later L2RPN competitions focused on trust and sending an alarm signal [26], and the inclusion of batteries [37].
Graph neural networks have been successfully applied to many tasks of power grid operation, among which power flow estimation [5, 4], optimal power flow computation [31, 11], load shedding [18], outage prediction [13], and stability prediction [30]. For a comprehensive survey on graph reinforcement learning for power grids, we recommend the survey by Hassouna et al. [14].
Several studies investigated GNNs for topology control. Yoon et al. [45] propose a hierarchical actor-critic approach that uses a GNN module to obtain graph embeddings. Their approach won the L2RPN WCCI 2020 challenge. Xu et al. [43] are the first to apply GNNs with graph attention layers for grid topology control. They apply this GNN with a proximal policy optimization approach with an action space searched with monte-carlo tree search. Their approach can, on average, maintain the system for four days in the 2020 NeurIPS L2RPN robustness track. Xu et al. [44] find that a deep reinforcement learning agent which uses features extracted with a GNN trains more quickly and achieve better performance than an agent without. Van der Sar et al. [35] propose a hierarchical actor-critic approach with one agent per substations. These agents use graph neural networks blocks, although the details are left unspecified. They apply this approach to the IEEE5-bus network. Qiu et al. [33] propose a method similar to Yoon et al.’s [45] and obtain similar results.
Taha et al. [40] propose a different approach than directly predicting topology actions with GNNs. They use GNNs for grid estimation. These estimates are then used in grid topology tree search. We will also discuss the approach by Zhao et al. [47], which is relevant because of their focus on generalization. They propose a proximal policy optimization for redispatching that relies on graph embeddings. The graph embeddings are generated by a GraphSAGE model trained with a unsupervised learning task. They find that this graph module can succesfully represent out-of-distribution grid changes. Other 2020 NeurIPS L2RPN submissions also used graph neural networks but were not notably successful [27].
Other work did not use graph neural networks. These approaches have mostly used reinforcement learning (RL). Chauhan et al. [2] propose the PowRL framework, which integrates a novel heuristic with an RL approach. They achieved the best results on the legacy 2020 L2RPN robustness track yet. Dorfer et al. [6] adapt the AlphaZero framework [38] for the power grid domain. They achieved top results in the legacy 2020 WCCI L2RPN competition. Pan et al. [32] investigate the vulnerability of RL methods to adversarial attacks. They propose a more robust agent trained adversarially. Subramanian et al. [39] perform an in-depth investigation of a simple agent on a single scenario. Manczak et al. [24] propose a hierarchical RL approach that separates substation and configuration selection.
Imitation learning (IL) has also been applied to the problem of grid topology control. De Jong et al. [15] recently performed the first investigation in IL as a stand-alone method for topology control. They found that both IL and hybrid agents could obtain good performance and inference speed combinations. Prior work on IL used IL as a ’warm start’ for RL. Lan et al. [20] first proposed IL to pre-train an RL model for grid control. Binbinchen [8] achieved second place in the 2020 NeurIPS L2RPN submission by expanding this approach. Lehna et al. [22, 21] refined this approach. They introduced topology reversal and a higher-performing N-1 expert agent [22], and a heuristic target topology approach [21]. The winning submission of the 2023 L2RPN competition also included this approach [19].
Other approaches used neither RL nor IL. Zhou et al. [48] use evolutionary strategies to train a model to narrow down the action space. The submission that won both 2020 NeurIPS L2RPN tracks used a similar approach [27]. Although not ML, the expert system developed by Marot et al. has also achieved notoriety [25].
@@ 3.1 Power Grid
In this study, we experiment on the IEEE14-bus system, displayed in Figure 1. We simulate this system with the rte_case14_realistic environment from the Grid2Op library [^3] This environment has since been deprecated. Future research should use environment l2rpn_case14_sandbox.. The power grid includes fourteen substations, five generators, 11 loads, and 20 power lines. One generator is solar, one nuclear, one wind-based, and two are thermal. The grid is divided into two sides. The high-voltage transmission side contains substations 0 to 4. The low-voltage distribution side contains substations 5 to 13. Lines 15 to 19 model the transformers connecting the two sides of the grid. We adjust the thermal limits to the values specified by Subramanian et al. [39]. This makes the transmission and distribution differences more pronounced and realistic.
We also investigate network variations with single lines disabled. We refer to these as N-1 networks. We use the N-1 networks to investigate the ability of agents to operate with line outages and to test the generalization of ML models to out-of-distribution networks.
Figure 1: The default state of Grid2Op environment rte_case14_realistic.
@@ 3.2 Operational Period
The environment features one thousand scenarios, each consisting of 8064 five-minute timesteps. One scenario thus models 28 days. Each scenario features different injection profiles. We split the scenarios into individual days. This daily period better reflects operational periods and emulates beneficial topology reversal [22]. A game-over occurs when the power grid fails to transport sufficient power from generators to loads [^4] Three scenarios were excluded where Grid2Op failed to converge the power flow without agent misoperation. These scenarios were omitted throughout the study..
@@ 3.3 Action Space
The substations have precisely two busbars in the environment. The topology actions switch objects, i.e., generators, loads, and line endpoints, between these busbars. The topology vector specifies the busbar attachments of objects. Each index in the vector corresponds to an object. At an index, a value of 1 indicates attachment to the first busbar, 2 indicates attachment to the second busbar, and -1 indicates object disconnection.
Actions are limited to a single substation per timestep to reflect real-world timing constraints. Certain substation configurations are invalid (isolated generators or loads) or redundant (configurations mirrored w.r.t. busbars). We use the approach by Subramanian et al. [39] to filter the space of configuration and set-actions. This produces the action space. The procedure is repeated for the different networks to generate the corresponding action spaces (see Subsection 4.1).

### v3

# Graph Neural Networks for Transmission Grid Topology Control: Busbar Information Asymmetry and Heterogeneous Representations

@@ Abstract
Factors such as the proliferation of renewable energy and electrification contribute to grid congestion as a pressing problem. Topology control is an appealing method for relieving congestion, but traditional approaches for topology discovery have proven too slow for practical application. Recent research has focused on machine learning (ML) as an efficient alternative. Graph neural networks (GNNs) are particularly well-suited for topology control applications due to their ability to model the graph structure of power grids. This study investigates the effect of the graph representation on GNN effectiveness for topology control. We identify the busbar information asymmetry problem inherent to the popular homogeneous graph representation. We propose a heterogeneous graph representation that resolves this problem. We apply GNNs with both representations and a fully connected neural network (FCNN) baseline on an imitation learning task. The models are evaluated by classification accuracy and grid operation ability. We find that heterogeneous GNNs perform best on in-distribution network configurations, followed by FCNNs, and lastly, homogeneous GNNs. We also find that both GNN types generalize better to out-of-distribution network configurations than FCNNs.
@@
Busbar information asymmetry limits graph neural networks for grid topology control.
Heterogeneous graph models of the grid increase model expressiveness and performance.
Machine learning agents obtain good grid operation performance at high efficiency.
Graph neural networks generalize better to out-of-distribution grid states.
@@ 1 Introduction
The energy transition, while critical, introduces significant new challenges to transmission system operators (TSOs) entseo_energy_transistion . The inherent variability and limited dispatchability of renewable energy sources can lead to supply peaks. TenneT, the Dutch TSO, projects an increase in installed renewable power capacity in the Netherlands from 29.6 GW in 2022 to 103.8 GW by 2033 tennet_2024 . Simultaneously, the electrification of sectors such as transportation is driving an increase in total electrical demand. TenneT forecasts Dutch electrical demand to increase from 113 TWh in 2022 to 171 TWh in 3033, representing a growth of over 50% tennet_2024 . Similar trends are anticipated in neighboring countries tennet_2024 . These developments introduce challenges such as grid congestion.
Grid congestion occurs when grid components approach or exceed their operational capacity limits. Congestion already poses a problem with tangible consequences: the European Union Agency for the Cooperation of Energy Regulators estimated the cost of European grid congestion in 2023 to be approximately 4 billion EUR ACER2024_CrossZonalCapacities . Furthermore, potential grid congestion scenarios can lead to the denial of connection requests for prospective customers, halting the development of businesses, schools, housing, and other facilities. These problems will quickly escalate without intervention.
Beyond the development of new infrastructure and the rejection of connection requests, TSOs can employ other measures. TSOs typically use redispatching and/or curtailment, involving the rebalancing or reduction of power production. However, both approaches are costly and limited in their applicability. Topology control presents an alternative method for congestion management that is cost-effective and flexible. Network topology can be altered by reconfiguring component connections within substations and by disabling power lines. The resulting change in power flow can alleviate congestion. Topology control thus present a cheap and flexible congestion management technique kelly2020rlelec .
Nevertheless, the usage of topology control by TSOs remains minimal. The number of possible network topologies grows exponentially with grid size, and most topologies cannot be operated. Moreover, topological changes affect power flow in complex, nonlinear ways that are time-consuming to compute. Finding the correct topology for a particular situation is thus difficult, and usage of topology control has hence remained minimal. However recent advancements have been made with the introduction of GridOptions by TenneT, a pioneering real-world tool for topological strategy suggestion jan_viebahn_2024 . The tool has demonstrated the value of topology control, showing that even simple topological strategies could reduce line congestion by 10 - 20%. The tool currently relies on fast load-flow calculations for screening topologies. However, the tool’s scope remains limited to a small section of the Dutch grid, including only eight substations, and the long inference time limits it to day-ahead application. To remove these limitations, the developers propose the future use of machine learning to speed up computation.
The idea of mitigating grid congestion with topology control dates back to the 1980s 4334990 . The discrete, nonlinear nature of topological actions limits the applicability of direct optimization methods and early research therefore often relied on computationally intensive simulations of topological actions. To address time constraints, simplified models or solvers such as linearized DC approximations were frequently used 515208 , 6939408 . Advanced methods have used more sophisticated solvers 1525118 or techniques to narrow the set of topological actions considered 7792703 . Nonetheless, these approaches often rely on inaccurate approximations, and their reliance on simulation renders them infeasible for large-scale real-world usage. Other research has focused on mathematical optimization techniques, such as mixed-integer programming (MIP) to approximate optimal busbar configurations. However, such approaches also prove too slow, with proposed approaches being limited to only line-switching 4492805 or busbar-switching within a single substation 9483070 .
Recently, machine learning (ML) techniques have garnered significant interest within the power systems community. ML models are notoriously good at learning complex nonlinear patterns, such as the patterns caused by topological actions. Furthermore, inference with ML models is quick as it does not require extensive search or optimization. The Learning to Run a Power Network (L2RPN) competitions marot2019learning , marot2021learning and the Grid2Op Python library grid2op  [^1] https://grid2op.readthedocs.io/ played a large part in the popularization of this research area. The field now has an active research community.
Power grid operation is ultimately a sequential decision-making problem. ML research on topology control has largely focused on the reinforcement learning Subramanian_2021 , 10.1609/aaai.v37i12.26724 and imitation learning lan2019aibased , lehna2023managing paradigms. Reinforcercement learning (RL) involves training agents by directly interacting with the environment. However, RL models for topology control often converge to “do-nothing” policies because the default topology is mostly satisfactory and topologies that improve upon it are rare githubGitHubAsprinChinaL2RPN_NIPS_2020_a_PPO_Solution . Imitation learning (IL) involves training a ML model to replicate the behavior of an expert agent. It can be used either as pre-training for RL to avoid the aforementioned problem or as a stand alone method deJong2024 . Evolutionary strategies have been also been proposed for the problem zhou2021actionsetbasedpolicy .
Graph neural networks (GNNs) 4700287 have recently recieved much attention, both from the machine learning and power engineering community. Deep learning models have traditionally been designed for tabular, sequential, or grid-like data. Consequently, ML methods for power grid operation have historically not or minimally exploited the network structure of power grids. This is a major limitation as power grid topologies are dynamic and frequently change, for example due to maintenance. Graph neural networks encode network structure into their representation and are therefore well-suited for power grid modeling. GNNs have hence been applied to various tasks related to grid operation, such as solving power flows 8851855 , DONON2020106547 , optimal power flow 9053140 , grid reliability or risk assessment CAMBIERVANNOOTEN2025125401 , ZHANG2025124793 , and state estimation NGO2024122602 . Similarly, GNNs have been applied to the problem of topology control with promising results yoon2020winning , 9830198 . However, despite the existing research, important topics have not yet been addressed by the literature:
@@
    Most present literature applying graph neural networks to topology control uses the same graph representation 9535409 , vandersar2023 , 9347305 , 9830198 , Hassouna2025LearningTA . In this representation, configurable objects are connected by edges if they are connected to the same busbar or power line. This is, however, only one of many ways to represent the power grid as a graph. On the related problem of optimal power flow, progress has been made by considering richer, heterogeneous graph representations ghamizi2024powerflowmultinet . However, for topology control, a study comparing or proposing improved graph representations for GNNs does not yet exist.
@@
    Power grids are frequently subject to changes, which cannot impede operation. Reliable and robust operation requires models to generalize or adapt to, potentially unexpected, altered network states. Prior work has focused on generalization to out-of-distribution injection profiles marot2021learning . However, no work has investigated generalization to out-of-distribution network configurations.
This study addresses these gaps in the literature. We identify a problem with the popular graph representation and propose a heterogeneous graph representation that resolves it. We train the GNNs, as well as ‘traditional’ fully connected neural networks (FCNNs), on a multi-label binary classification imitation learning task. The utilized dataset consists of various topologies and outages deJong2024 . The models are evaluated by their classification accuracy and their efficacy when used as agents to operate a simulated power grid. Furthermore, we compare the operation efficacy and inference speed of ML models against expert and hybrid agents. We also investigate the accuracy and operation efficacy of models to out-of-distribution network configurations. In particular, the foremost contributions of this study are:
@@
    We perform the first analysis of the effect of power grid graph representation on GNN topology control performance. The currently popular graph representation models only the current connections between objects. However, it ignores which objects might be connected by actions, thereby obscuring the effects of actions. We call this the ‘busbar information asymmetry’, which hinders information flow and limits GNN expressiveness. We propose a heterogeneous graph representation that models both current and potential connections and consequently resolves the asymmetry. We evaluate both graph representations with GNNs.
@@
    We perform the first investigation of the ability of graph neural networks to generalize topology control to out-of-distribution grid configurations. We evaluate this both by model accuracy and grid operation effectiveness.
This paper is organized as follows. Section 2 describes the power grid setup used in this study. Section 3 describes the methods used to develop the ML models. Section 4 describes the evaluation and analysis of the ML models. Sections 5 provides discussion and recommendations for future work, respectively. The code for this project is available online.  [^2] https://github.com/MatthijsdeJ/GNN_PN_Imitation_Learning Table 1 lists the meaning of mathematical symbols used in this paper.
@@
$O$ Set of grid objects (i.e., generators, loads, and line endpoints).
$S$ Set of substations.
$\textbf{x}_{u}$ Features of a GNN model at node $u\in O$.
$\textbf{h}_{u,k}$ Node embeddings of a GNN model at node $u\in O$ in layer $k$.
$\mathcal{N}(u)$ Neighbors of node $u\in O$.
$W$ Weight matrix.
$\sigma$ Activation function.
$\textbf{p}=(0,1)^{|O|}$ Output of a ML model.
$\textbf{y}=\{0,1\}^{|O|}$ Target of a ML model.
$\alpha=0.1$ Label weight hyperparameter.
$\textbf{w}=\{\alpha,1\}^{|O|}$ Label weights for the loss.
@@
$\{\textbf{p}_{s}\}{}\forall s\in S$ Partition of the output vector p into subvectors $\textbf{p}_{s}$. Subvector $\textbf{p}_{s}$ corresponds to substation $s\in S$.
$\eta$ = 0.97 Activity threshold parameter.
$\theta=1.0$ Risk threshold parameter, used by various agents.
@@ 2.1 Power Grid
The IEEE 14-bus system (Fig. 1) was used for the experiment in this study. Simulations were conducted using rte_case14_realistic environment of the Grid2Op library [^3] Note that this environment has since been deprecated. Future research should use environment l2rpn_case14_sandbox.. The system comprises 14 substations, five generators (one solar, one nuclear, one wind-based, and two thermal), 11 loads, and 20 power lines. Substations 0-4 represent the high-voltage transmission side, while substations 5-13 represent the low-voltage distribution side, interconnected by transformers modeled by lines 15-19. To amplify the differences between transmission and distribution characteristics, thermal limits were adjusted according to the values specified by Subramanian et al. Subramanian_2021 .
Furthermore, we investigate network variations/configurations with single lines disabled, referred to as N-1 networks. These N-1 networks were used to investigate the ability of agents to operate with line outage conditions and to assess the generalization of ML models to out-of-distribution network configurations.
Figure 1: The default state of Grid2Op environment rte_case14_realistic. Numbers indicate power line or substation IDs.
@@ 2.2 Operational Period
The environment features 1000 distinct scenarios, each spanning 8064 five-minute timesteps, thus modeling a 28-day period. Each scenario is characterized by unique injection profiles. We split the scenarios into individual days, creating 28,000 episodes. This day-long period better reflects operational periods, and resetting the topology between days emulates beneficial topology reversal lehna2023managing . A simulation terminates early (“game over”) when the power grid is unable to transport sufficient power from generators to loads [^4] In three scenarios, the Grid2Op power flow solver failed to converge, not as a result of agent misoperation. These scenarios were omitted throughout the study..
@@ 2.3 Action Space
Within the simulation environment, each substation is modeled with two busbars. Topology control actions involve switching network objects (generators, loads, and line endpoints) between these busbars. A topology vector specifies the busbar attachment of each object; each index in the vector corresponds to a specific object, with a value of 1 indicating connection to the first busbar, 2 indicating connection to the second busbar, and -1 indicating disconnection.
Actions are limited to a single substation per timestep, reflecting real-world operational constraints. Certain substation configurations are considered invalid (e.g., isolated generators or loads) or redundant (configurations mirrored across busbars). The action space is generated by applying the filtering approach developed by Subramanian et al. Subramanian_2021 . The procedure is repeated for each network variation to generate the corresponding action spaces (see Subsection 3.1).